# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-tuning

### v1

@@ 6.2 Experiment Results
In Table 1, we present a comprehensive evaluation of safety fine-tuned VLMs across three key metrics: safety performance against unsafe input queries (measured by ASR, where lower is better), over-prudence performance against safe input queries (measured by RR, where lower is better), and model utility on representative downstream tasks (measured by Acc, where higher is better). Recall that both ASR and RR are evaluated before and after exploiting spurious correlations, i.e., promoting non-rejection bias and rejection bias via one-word modification, respectively.
@@
We draw some key observations below. First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) exhibit a safety mirage, as evidenced by a significant rise in ASR after applying the one-word attack–nearly 60% on the VLGuard unsafe evaluation set and nearly 30% on the SPA-VL unsafe evaluation set of LLaVA-1.5-7B based model. Additionally, these baselines exhibit a significant over-prudence issue, as evidenced by an over 90% increase after one-word modification in RR on the safe evaluation sets in VLGuard and SPA-VL. Furthermore, both full and LoRA fine-tuning exhibit a similar safety mirage phenomenon in baselines.
Second, compared to baselines, the unlearning-based approaches (NPO and RMU) exhibit significantly lower ASR increases after the one-word attack and maintain a low RR against safe queries, effectively alleviating both jailbreaking susceptibility and over-prudence. Compared to NPO, RMU achieves slightly better performance in both ASR (lower vulnerability to the one-word attack) and RR (reduced over-prudence). This result is consistent with LLM unlearning, where RMU typically outperforms NPO in knowledge unlearning [26]. Furthermore, the advantages of MU persist in both full fine-tuning and LoRA fine-tuning.
Third, from the perspective of model utility evaluation, we observe that unlearning-based approaches lead to a slight decrease in Acc on downstream tasks (approximately 1%), compared to supervised safety fine-tuning. This suggests a potential tradeoff between unlearning effectiveness and utility preservation, indicating that erasing harmful knowledge in VLMs may have an impact on general performance. Further optimizing this tradeoff remains an important future research direction to enhance VLM unlearning.
Table 2: Analyzing the safety mechanisms of unlearning-based approaches vs. baselines. Before and after applying the one-shot “What”-initialized attack, the safety rate (1-ASR) against unsafe queries is decomposed into RR (rejection rate) and irrelevance rate (IR), where IR represents responses that are irrelevant to the unsafe queries. All other setups remain consistent with Table 1.
@@
Other safety evaluation. We further evaluate different safety fine-tuning approaches using LLaVA-v1.5-7B on MM-SafetyBench and FigStep, with results Table A1. Consistent with our findings on VLGuard and SPA-VL, the one-word attack significantly increases ASR across both benchmarks. In Fig. 5, we present input-output demonstrations of safety fine-tuned LLaVA-1.5-7B, comparing the baseline Mixed-SFT with the RMU-based unlearning approach. The original LLaVA-1.5-7B model is vulnerable to unsafe queries both before and after the one-word attack. Here the attack is primarily executed by replacing “How” with “What” in text queries. However, the one-word attack effectively bypass the safety mechanism from Mixed-SFT, triggering unsafe response against the query starting with “What”. Moreover, the Mixed-SFT fine-tuned VLM exhibits over-prudence, as evidenced by its rejection response to a safe text query starting with “Share”. Here “Share” is identified as a high-frequency word in the fine-tuning dataset that is strongly correlated with rejection responses in Fig. 3. By contrast, the RMU unlearning approach effectively defends against the one-word attack while also mitigating the over-rejection issue.
Figure 6: Input token sensitivity analysis for “What”-initiated unsafe queries and “Share”-initiated safe queries over VLGuard using LLaVA-1.5-7B Mixed-SFT, before and after masking “What” and “Share”. Sensitivity is measured using per-token masking (i.e., replacing the original token with a blank placeholder [PAD]) to evaluate each token’s influence on response rejection probabilities.
Irrelevance response for safety enhancement by unlearning. To understand the difference in the safety mechanism of MU-based approaches vs. safety-aware SFT, Table 2 presents the rates of unsafe, irrelevant, and rejection responses in our safety evaluation, both before and after a 1-shot one-word attack. Here, the safety rate (1-ASR) is further decomposed into: (1) irrelevant responses where the model sidesteps the unsafe query by generating a response that is unrelated to the harmful content, and (2) rejection responses where the model explicitly refuses to respond. As observed, conventional SFT-based safety fine-tuning strategies predominantly produce rejection-based responses, as reflected in the overly high RR in both “Before” and “After” scenarios. In contrast, our unlearning-based methods primarily yield irrelevant responses, reducing the model’s reliance on outright rejections, as evidenced by the high irrelevance rate (IR). Notably, the RMU-Unlearning approach steers the model’s output distribution closer to a random vector, making responses more likely to be classified as irrelevant rather than explicitly rejecting queries.
Input token saliency analysis. Recall that spurious correlations emerge between specific words in textual queries and safety labels. To further analyze this, we investigate input token sensitivity by applying per-token masking (i.e., replacing the original input token with a blank placeholder [PAD]) to assess the token influence in non-rejection and rejection responses. Fig. 6 presents the rejection probability using LLaVA-1.5-7B Mixed-SFT for all “What”-initiated unsafe input queries and “Share”-initiated safe queries before and after masking “What” and “Share”, respectively, over the VLGuard dataset. As we can see, masking “What” significantly increases the rejection probability, confirming its role in inducing non-rejection bias in VLGuard. Masking “Share” significantly reduces the rejection probability compared to its unmasked counterpart, demonstrating its influence in reinforcing rejection bias. See Appendix B.6 for specific textual input saliency maps.
Additional results. In Appendix B.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
@@ 7 Conclusion
In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that emerges from supervised safety fine-tuning. Our analysis reveals that biases in the fine-tuning dataset reinforce spurious correlations between superficial textual patterns and safety labels, leading to a false sense of security. As a result, fine-tuned VLMs become highly susceptible to simple one-word jailbreaking attacks while also exhibiting over-prudence, unnecessarily rejecting benign queries. To address these, we propose MU as a principled alternative to fine-tuning. Unlike supervised approaches, MU directly removes harmful knowledge without relying on explicit safety labels, thereby avoiding biased feature-label mappings and mitigating spurious correlations. Extensive experiments confirm the existence of the “safety mirage” in conventional VLM safety fine-tuning and demonstrate that MU-based safety alignment significantly enhances robustness against jailbreaking attacks, reduces over-prudence, and preserves strong performance on standard VQA tasks.
8 Broader Impacts
Our work reveals spurious correlations in VLM safety fine-tuning, which are prevalent across models and can function as unintentional “backdoors” in real-world applications. Malicious actors could exploit these correlations to launch jailbreaking attacks, potentially extracting sensitive or private information from VLMs. While the advocated unlearning methods aim to enhance safety alignment, they also present dual-use risks. A bad actor could misuse these techniques to erase safety guardrails, rendering a previously robust model unsafe. If such an unlearned model were publicly released, it could bypass safety mechanisms, increasing the risk of harmful content generation.
@@ 9 Acknowledgement
We extend our gratitude to Cisco Research and DSO National Laboratories for their support in this project. Additionally, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the National Science Foundation (NSF) CISE Core Program Award IIS-2207052, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the NSF CAREER Award IIS-2338068, the Amazon Research Award, and the IBM PhD Fellowship.
@@
A One-word Over Prudence
Similar to the one-word jailbreaking attack, inserting a rejection bias-inducing word (e.g., “Share” as shown in Fig. 3) into text queries causes the fine-tuned VLM to produce no meaningful output, even for benign inputs. We also apply a multi-shot strategy — modifying different paraphrases of a benign query with “Share”. Notably, even a one-shot “Share” modification leads to a 90% over-rejection rate on safe queries.
@@ A One-word Over Prudence
Figure A1: Rejection rate vs. $K$, evaluated before and after applying the “Share”-initialized one-word modification to safe input queries, cause over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model [23].
B Additional Experiment Details and Results
B.1 Additional Fine-tuning Setups
All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs. We perform fine-tuning using both full-parameter and LoRA-based methods. For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For LoRA-based fine-tuning, following [6, 21], we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$. In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
B.2 Machine Unlearning Implementation Details
@@
To address problem (2), fine-tuning can be effectively performed either through full parameter updates or parameter-efficient techniques such as low-rank adaptation (LoRA) [82]. Common implementations adopt cross-entropy loss for $\ell_{\mathrm{r}}$, while $\ell_{\mathrm{u}}$ can flexibly incorporate various optimization objectives such as cross-entropy loss, DPO [76], PPO [83].
@@
To apply the RMU algorithm for safety alignment, we use unsafe input-output pairs from VLGuard as the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ for unlearning. We employ Llama-2-13B-Chat to verify the harmfulness of the original LLaVA series model’s responses to unsafe queries and select the most harmful response as the target for unlearning. Specifically, for RMU, we combine the question and answer into a single instance as VLM’s input to be forgotten. For the retain dataset $\mathcal{D}_{\mathrm{r}}$, we use a mixture of the LLaVA fine-tuning data and the safe input-output pairs from VLGuard. The unlearning loss $\ell_{\mathrm{u}}$ defined as (3), where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations (we use layers 4, 5, and 6 in our implementation), $c$ is a hyperparameter for activation scaling (set to $c=10$), and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.2$.
@@
Compared to the RMU implementation, we use the same settings for the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ and the retain dataset $\mathcal{D}_{\mathrm{r}}$. However, for NPO, the input to the VLM from $\mathcal{D}_{\mathrm{u}}$ consists solely of the harmful question. In this setting, unsafe data designated for unlearning is treated as negative examples within a direct preference optimization framework [76]. The unlearning training objective is defined as (5), where $\sigma(\cdot)$ denotes the sigmoid function, $\beta>0$ is a temperature parameter (set to $\beta=0.6$ in our implementation), $\pi_{{\bm{\theta}}}(\mathbf{x})$ is the prediction probability of the model given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}(\mathbf{x})$ is the corresponding probability from the reference model (i.e., the initial model prior to unlearning). The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.0$.
B.3 Additional Baseline Details
For the Unsafe-Filter baseline, we employ the LLaMA-Guard-3-11B-Vision [81] model to assess the safety of image-instruction-answer pairs in the LLaVA pre-training and fine-tuning datasets. Unsafe data is filtered out, and LLaVA is subsequently fine-tuned using only the remaining safe samples. The fine-tuning setup follows the original LLaVA configuration.
B.4 Evaluation Setup Details
To validate the harmfulness of the responses generated by the fine-tuned model, we employ =Qwen2.5-VL-7B-Instruct model as a judge. Below is the prompt used for evaluating content relevance, which is included in our experiments:

### v2

@@ 6.2 Experiment Results
Table 1 summarizes results across three metrics: ASR (safety), RR (over-prudence), and Acc (utility). Both ASR and RR are evaluated before and after exploiting spurious correlations, i.e., promoting non-rejection bias and rejection bias via one-word attack and one-word modification. We draw some key observations below.
First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) suffer from a clear safety mirage: ASR rises sharply after one-word attacks (nearly 60% on VLGuard-unsafe and 30% on SPA-VL-unsafe for 7B models), while RR also increases dramatically (over 90%) on safe queries. This vulnerability appears consistently under both full and LoRA fine-tuning. Second, unlearning-based approaches (NPO and RMU) substantially reduce ASR increases and maintain low RR, alleviating both jailbreaking susceptibility and over-prudence. Compared to NPO, RMU performs slightly better. Third, in terms of general utility, unlearning-based methods (NPO and RMU) show only a slight accuracy drop (around 1%) compared to the original model. This is expected, as a trade-off between robustness and utility exists, and the gain in safety robustness far outweighs the minor utility loss using unlearning.
To better balance safety and utility in VLMs, we can adopt the idea of coreset unlearning (Pal et al., 2025). Using fewer safety samples for unlearning-based fine-tuning improves utility while still achieving stronger safety performance than conventional fine-tuning baselines. Detailed results are provided in Table A1 of Appendix D.4.
Other safety evaluation. We further evaluate different safety fine-tuning approaches using LLaVA-v1.5-7B on MM-SafetyBench and FigStep, with results Table A4 of Appendix D.5. Consistent with our findings on VLGuard and SPA-VL, one-word perturbations remain highly effective, and MU-based methods are more effective and robust than conventional safety fine-tuning.
In Fig. 6, we show input-output examples of safety fine-tuned LLaVA-1.5-7B, comparing Mixed-SFT with RMU-based unlearning. The original model remains vulnerable to unsafe queries both before and after the one-word attack, implemented by replacing “How” with “What.” While Mixed-SFT reduces some unsafe outputs, the attack still bypasses its safety filter and triggers unsafe responses. Mixed-SFT also suffers from over-prudence, rejecting a benign query starting with “Share,” which is spuriously correlated with rejection responses in Fig. 3. In contrast, RMU-based unlearning both resists the one-word attack and alleviates over-rejection on safe queries.
@@
Unlearning as a distinct safety mechanism: A response analysis. To highlight the difference in safety mechanisms between MU-based approaches and safety-aware SFT, Table 2 reports the proportions of unsafe, irrelevant, and rejection responses before and after 1-shot one-word attack. Here, the safety rate (1–ASR) is decomposed into: (1) irrelevant responses, where the model sidesteps the unsafe query with unrelated content, and (2) rejection responses, where the model explicitly refuses to answer. Conventional SFT approaches, guided by safety labels, rely mainly on rejection responses, yielding consistently high RR both “Before” and “After”. In contrast, MU-based methods do not rely on explicit rejection labels. Instead, they achieve safety primarily through irrelevant responses, reflecting a more label-free mechanism for mitigating spurious correlation.
Table 2: Analyzing the safety mechanisms of unlearning-based approaches versus baselines. ASR was reported both before and after the 1-shot “what” initialized attack. Safety rate is decomposed into the rejection rate ($R$) and irrelevance rate (IR), where IR denotes non-rejection outputs judged as irrelevant to the unsafe queries. All other setups remain consistent with Table 1.
@@
Figure 7: Input token sensitivity analysis for “What”-initiated unsafe queries and “Share”-initiated safe queries over VLGuard using LLaVA-1.5-7B Mixed-SFT, before and after masking “What” and “Share”. Sensitivity is measured using per-token masking (i.e., replacing the original token with a blank placeholder [PAD]) to evaluate each token’s influence on rejection probabilities.
Validating spurious correlations via input token saliency analysis. We further validate spurious correlations between specific query words and safety labels using token-level sensitivity analysis, where selected input tokens are masked (i.e., replaced with [PAD]) and their effect on rejection probabilities is measured. Figure 7 shows the results for “What’ initiated unsafe queries and “Share” initiated safe queries using LLaVA-1.5-7B Mixed-SFT. Masking “What” leads to a sharp increase in rejection probability, confirming its role in inducing non-rejection bias. Conversely, masking “Share” reduces the rejection probability compared to the unmasked case, highlighting its influence in reinforcing rejection bias. Other saliency maps are provided in Appendix D.6.
Additional results. In Appendix D.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
@@ 7 Conclusion
In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that arises from supervised safety fine-tuning. We show that biases in fine-tuning datasets reinforce spurious correlations between superficial textual patterns and safety labels, creating a false sense of security. As a result, fine-tuned VLMs remain vulnerable to simple one-word jailbreaking attacks and exhibit over-prudence, unnecessarily rejecting benign queries. To address this, we propose machine unlearning (MU) as a principled alternative: rather than relying on explicit safety labels, MU removes harmful knowledge in a label-free manner, mitigating spurious correlations. Our experiments validate the safety mirage phenomenon and demonstrate that MU improves robustness against jailbreaks, alleviates over-prudence, and preserves utility on standard VQA tasks. We refer readers to Appendix FE for limitations, broad impacts, and LLM usage.
@@ 8 Acknowledgement
The project is primarily supported by research awards from Cisco Research and DSO National Laboratories. In addition, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the NSF CISE Core Program Awards IIS-2207052 and IIS-2504263, the NSF CAREER Award IIS-2338068, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the Amazon Research Award, the Open Philanthropy Research Award, and the IBM PhD Fellowship.
@@
Appendix A Visualization of Motivation Samples
Figure A1: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-SPA-VL-DPO-30K and LLaVA-v1.5-7B-SPA-VL-PPO-30K (Zhang et al., 2024a; Rafailov et al., 2023; Schulman et al., 2017)). The visual marks keep consistent meaning with Fig. 2. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. But if the questions use the starting word “Please”, it would bypass this safeguard. (b) Over-prudence: The fine-tuned model will generate the benign responses when the queries are both harmless, but ask the problems starting from “What” can trigger unnecessary refusals.
Appendix B Additional Results on Spurious Correlations
B.1 Additional Results of One-Word Jailbreaking
Figure A2: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “Please”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL unsafe test set.
For the SPA-VL test set, we further evaluated the LLaVA-v1.5-7B-PPO-30K model using the same $K$-shot paraphrasing protocol as in the main text. Here, we selected “please” as the adversarial biasing word for unsafe queries. The motivation is dataset-driven: in the SPA-VL training corpus, the token “please” occurs with very low frequency like Fig. 3 shows, so asserting it to unsafe queries disrupts the model’s learned phrasing patterns. Under PPO-based preference optimization, this rare starter token reduces the model’s tendency to follow the preferred safe/rejection response, thereby substantially increasing the attack success rate (ASR). These results corroborate our main finding that low-frequency tokens can serve as effective adversarial triggers in one-word jailbreaking. Quantitatively, as shown in Fig. A2, the ASR rises from about 12% before modification to nearly 30% at $K=1$, and continues to grow steadily with larger $K$, ultimately exceeding 60%. This confirms that the “Please” based attack is consistently effective across different shot settings.
B.2 Additional Results of One-Word Over-prudence
@@
Figure A3: Rejection rate of $K$-shot one-word modification, evaluated before and after applying the “What” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL safe.
We also tested the over-prudence effect on SPA-VL safe queries. For SPA-VL, we used “what” as the biased starter word. Unlike “please”, “what” is highly frequent in the SPA-VL training corpus as Fig. 3 presents. When combined with PPO preference optimization, this frequent token strongly biases the model toward the preferred rejection-style responses. As a result, benign queries initialized with “what” exhibit a substantially increased rejection rate (RR), confirming the over-prudence phenomenon. These findings highlight how high-frequency tokens can act as rejection-bias triggers in safe contexts, complementing the low-frequency jailbreak triggers observed in unsafe contexts. From Fig. A3, we see that the RR almost doubles, jumping from about 39% before modification to nearly 79% after just one shot, and quickly saturating around 90% as $K$ grows. This illustrates the strong over-rejection tendency induced by the “what” starter.
Appendix C Implementations Details of Machine Unlearning
@@
To address problem (2), fine-tuning can be effectively performed either through full parameter updates or parameter-efficient techniques such as low-rank adaptation (LoRA). Common implementations adopt cross-entropy loss for $\ell_{\mathrm{r}}$, while $\ell_{\mathrm{u}}$ can flexibly incorporate various optimization objectives such as cross-entropy loss, DPO (Rafailov et al., 2023), PPO (Schulman et al., 2017).
Details of Retain Loss $\ell_{\mathrm{r}}$.
To address the problem, directly applying MU objectives to VLMs often leads to instability or even model collapse, so we design $\ell_{\mathrm{r}}$ as a composite of two terms in (6). For the $\ell_{\mathrm{ft}}$, it is the common implementations that adopt cross-entropy loss on $\mathcal{D}_{\mathrm{r}}$ like the original LLaVA fine-tune. While for the $\ell_{\mathrm{mu},\mathrm{r}}$, it is depend on the choices of $\ell_{\mathrm{u}}$, we set it as the corresponding retain loss of the specific unlearning algorithm. For example, when $\ell_{\mathrm{u}}$ is the RMU training loss,
$\ell_{\mathrm{\mathrm{mu},\mathrm{r}}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{r}}}\big[\|M_{{\bm{\theta}}}(\mathbf{x})-M_{{\bm{\theta}}_{\mathrm{ref}}}(\mathbf{x})\|_{2}^{2}\big],$ (A1)
where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations of $\bm{\theta}$}, and $M_{{\bm{\theta}}_{\mathrm{ref}}}(\cdot)$ those of the reference (pre-unlearning) model. While for $\ell_{\mathrm{u}}$ equals to the NPO training loss,
$\ell_{\mathrm{\mathrm{mu},\mathrm{r}}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{r}}}\left[-\frac{2}{\beta}\log\sigma\left(\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],$ (A2)
where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter, $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the current model given the retain input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning. During the experiment, we both set $\alpha=1$ in (2).
@@
To apply the RMU algorithm for safety alignment, we use unsafe input-output pairs from VLGuard as the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ for unlearning. We employ Llama-2-13B-Chat to verify the harmfulness of the original LLaVA series model’s responses to unsafe queries and select the most harmful response as the target for unlearning. Specifically, for RMU, we combine the question and answer into a single instance as VLM’s input to be forgotten. For the retain dataset $\mathcal{D}_{\mathrm{r}}$, we use a mixture of the LLaVA fine-tuning data and the safe input-output pairs from VLGuard. The unlearning loss $\ell_{\mathrm{u}}$ defined as (3), where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations (we use layers 4, 5, and 6 in our implementation), $c$ is a hyperparameter for activation scaling (set to $c=10$), and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. The retain loss $\ell_{\mathrm{r}}$ is computed data, with the regularization parameter set to $\gamma=1.2$.
@@
Compared to the RMU implementation, we use the same settings for the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ and the retain dataset $\mathcal{D}_{\mathrm{r}}$. However, for NPO, the input to the VLM from $\mathcal{D}_{\mathrm{u}}$ consists solely of the harmful question. In this setting, unsafe data designated for unlearning is treated as negative examples within a direct preference optimization framework (Rafailov et al., 2023). The unlearning training objective is defined as (5), where $\sigma(\cdot)$ denotes the sigmoid function, $\beta>0$ is a temperature parameter (set to $\beta=0.6$ in our implementation), $\pi_{{\bm{\theta}}}(\mathbf{x})$ is the prediction probability of the model given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}(\mathbf{x})$ is the corresponding probability from the reference model (i.e., the initial model prior to unlearning). The retain loss $\ell_{\mathrm{r}}$ computed with the regularization parameter set to $\gamma=1.0$.
Appendix D Additional Experiment Details and Results
D.1 Additional Fine-tuning Setups
All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs. We perform fine-tuning using both full-parameter and LoRA-based methods. For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For LoRA-based fine-tuning, following (Liu et al., 2023; Zhang et al., 2024a), we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$. In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
D.2 Additional Baseline Details
An additional SFT baseline is constructed by filtering out unsafe samples at both stages of LLaVA training. Specifically, we employ LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to assess the safety of all image–instruction–answer triplets in LLaVA’s pre-training and instruction-tuning datasets. Any sample flagged as unsafe is removed prior to training. As a result, the model is pre-trained exclusively on safe image–text pairs, and subsequently fine-tuned on safe instruction–answer pairs, ensuring that unsafe data is excluded from the entire training pipeline. The fine-tuning setup otherwise follows the original LLaVA configuration, making this a direct comparison point for evaluating the effect of safety filtering throughout both training stages.
D.3 Evaluation Setup Details
To validate the harmfulness of the responses generated by the fine-tuned model, we employ Qwen2.5-VL-7B-Instruct model as a judge. Below is the prompt used for evaluating content relevance, which is included in our experiments: